PLease answer all queations ASAP! wil rate

1.

Georgia, Inc. has collected the following data on one of its products. The actual cost of the direct materials used is:

Direct materials standard (3 lbs @ $2/lb) $6 per finished unit
Total direct materials cost variance—unfavorable $14,250
Actual direct materials used 100,000 lbs.
Actual finished units produced 25,000 units

Multiple Choice

  • $135,750.

  • $164,250.

  • $150,000.

  • $182,000.

  • $100,000.

  • 2.

    Use the following data to find the total direct labor cost variance if the company produced 3,500 units during the period.

    Direct labor standard (4.00 hrs. @ $7.10/hr.) $ 28.40 per unit
    Actual hours worked 11,900
    Actual rate per hour $ 7.70

    Multiple Choice

  • $8,400 unfavorable.

  • $7,770 favorable.

  • $11,900 favorable.

  • $7,770 unfavorable.

  • $8,400 favorable

  • 3.

    Grant Co. uses the following standard to produce a single unit of its product: Variable overhead (1.60 hrs. per unit @ $3.70/hr.) Actual data for the month show total variable overhead costs of $138,640, and 22,000 units produced. The total variable overhead variance is:

    Multiple Choice

  • $8,400F.

  • $8,400U.

  • $62,400U.

  • $62,400F.

  • $0.

  • 4.

    The following information relating to a company's overhead costs is available.

    Actual total variable overhead

    $ 100,000

    Actual total fixed overhead

    $ 19,000

    Budgeted variable overhead rate per machine hour

    $ 3.00

    Budgeted total fixed overhead

    $ 15,000

    Budgeted machine hours allowed for actual output

    37,000


    Based on this information, the total variable overhead variance is:

    Multiple Choice

  • $11,000 favorable.

  • $5,000 favorable.

  • $11,000 unfavorable.

  • $5,000 unfavorable.

  • $4,459 favorable

  • 5.

    Sanchez Company's output for the current period was assigned a $434,000 standard direct labor cost. The direct labor variances included a $10,850 unfavorable direct labor rate variance and a $4,340 favorable direct labor efficiency variance. What is the actual total direct labor cost for the current period?

    Multiple Choice

  • $418,810.

  • $449,190.

  • $444,850.

  • $440,510.

  • $427,490.
Answer #1

Answer 1-

Total unit produced =25000

Standard cost/unit =$6

Total standard cost= 150000

Add unfavorable variance = 14250

Actual cost of direct material used= 164250

Answer 2-

Actual direct labor cost=

Actual Hours worked×Actual rate per hour

= 11900×7.7= 91630

Standard direct labor= 28.40×3500 =99400

If actual is less than standard than variance is favorable

Hence 99400-91630= 7700 favorable

Answer 3-

Standard per unit variable overhead= 1.6×3.7=5.92

Total production= 22000,

Standard variable overhead = 5.92×22000=130240

Actual variable overhead= 138640

Standard- Actual

= 130240-138640

=8400 Unfavorable

Answer 4-

Total standard variable cost = Budgeted hour rate × Budgeted hours

= 3 × 37000 = 111000

Actual variable cost = 100000

Total variable overhead variance = Budgeted - Actual

=111000 - 100000

=11000 favorable

Answer 5-

Standard direct labour cost = 434000

Less Direct labour variances = 10850

Add Direct labour rate variance = 4340

Actual total direct labour cost = 434000 - 10850 + 4340

= 427490

Note: All the answers are in Dollar($) only.